OpenStudy (loser66):

Find the radius of converges of \[\sum_{n=1}^\infty n^nx^{n^2}\] Please, help

OpenStudy (loser66):

@amistre64

OpenStudy (anonymous):

Have you tried the ratio or root tests?

OpenStudy (loser66):

I did Hadamard test.

OpenStudy (loser66):

But not sure about \(x^{n^2}\)

OpenStudy (loser66):

Is it \(x^{n^2}=x^{2n}\)?

OpenStudy (anonymous):

I had to look it up, but it strongly resembles the traditional root test. Setting up the limit, you have \[\lim_{n\to\infty}\left|n^nx^{n^2}\right|^{1/n}=\lim_{n\to\infty} n \left|x^n\right|\] since \(\left(n^n\right)^{1/n}=n\) and \(\left(x^{n^2}\right)^{1/n}=x^{n^2/n}=x^{n}\). The series converges for \(x\) that make the limit less than \(1\). What's the value of the limit?

OpenStudy (loser66):

Is it infinitive?

OpenStudy (anonymous):

Yes, the limit is infinity. What does this tell you about the series?

OpenStudy (loser66):

But if I apply Hadamard with \(\lim_{n\rightarrow \infty }|\sqrt[n]{n^n}|\) . What I am not sure is about n . On lecture notes, I have \(lim_{n\rightarrow \infty }\sqrt[n]{n^k}=lim_{n\rightarrow \infty}(\sqrt[n]{n})^k=1^k=1\)

OpenStudy (loser66):

So that if n =k (why not? right?) hence I still have lim =1 and then the sum converges and the radius is 1.

OpenStudy (loser66):

However, if counting like \(\sqrt[n]{n^n}=n\), then I got lim goes to infinitive and the sum is diverge. They conflict, right?

OpenStudy (anonymous):

I don't think you can simply exchange \(n\) with \(k\) in the work on the board. It looks like \(k\) is fixed, but \(n\) is not (because it is constant increasing, i.e. \(n\to\infty\)).

OpenStudy (anonymous):

So it's okay to say that \(\sqrt[n]{n^2}\to1^2\), but the reasoning does not hold if the exponent is not fixed.

OpenStudy (loser66):

Got you, so that, in this case, we have the series divergence and the radius is infinitive, right?

OpenStudy (anonymous):

Right.

OpenStudy (loser66):

Thanks for making it clear. I do appreciate.

OpenStudy (loser66):

This is the way my prof solved the problem: I post to let remind later, \[\sum_{n=1}^\infty n^n x^{n^2}=\sum_{k=1}^\infty a_kx^k\] \[a_k=\begin{cases}n^n~~~if ~~~k=n^2\\0~~~otherwise\end{cases}\]

OpenStudy (loser66):

\[\sqrt[k]a_k=\begin{cases}\sqrt[n^2]n^n =n^{1/n}~~~if~~~k=n^2\\0~~~~otherwise\end{cases}\]

OpenStudy (loser66):

As \(n\rightarrow \infty\), \(n^{1/n}\rightarrow 1\) hence \(lim_{k\rightarrow \infty}sup\sqrt[k]{a_k} =1\) That shows Radius \(R=\dfrac{1}{lim_{k\rightarrow \infty}\sqrt[k]{|a_k|}}=1/1=1\)

OpenStudy (anonymous):

Terribly sorry, I was mistaken. Apparently, the root (Hadamard) test applies for series strictly of the form \(\sum a_n x^n\), not \(\sum a_n x^{n^2}\). The latter is not a power series.